OpenStudy (anonymous):

Find the vertex of the function y = -x2 + x

OpenStudy (lieutenantgeneral):

www.wolframalpha.com

OpenStudy (tkhunny):

Your task is to Complete the Square! y = -x^2 + x y = -(x^2 - x + _____) + _____ (-1/2)^2 = 1/4 y = -(x^2 - x + 1/4) + 1/4 y = -(x - 1/4)^2 + 1/4 or y - 1/4 = -(x - 1/4)^2 The Vertex is now rather staring at you. That's why we call it Vertex Form!